### v1

# An Equation of State for Turbulence in the Gross-Pitaevskii model

Abstract
We report the numerical observation of a far-from-equilibrium equation of state (EoS) in the Gross-Pitaevskii model. This universal dimensionless EoS is constructed by relating the turbulent cascade’s scale-free spectrum amplitude to the energy flux of the steady state that emerges from the large-length-scale driving of a Bose-Einstein condensate (BEC). Remarkably, this EoS defies the generic predictions of wave-turbulent kinetic theory, even though the cascade spectra are quantitatively well understood within that theory. Finally, we find that the concept of quasi-static thermodynamic processes between equilibrium states extends to far-from-equilibrium steady states.
Equilibrium and near-equilibrium thermodynamics are conceptual cornerstones in physics, reducing the description of complex many-body systems of many degrees of freedom to relationships between a few macroscopic observables. When a system is in a far-from-equilibrium state, such a general framework does not exist, and unveiling universal features in such scenarios has been a major goal in modern physics. In situations where local equilibrium holds, the framework of local thermodynamics and hydrodynamics can provide a bridge to universal descriptions [1, 2, 3]. In the cases where even microscopic equilibrium is not realized, universal features have also been predicted theoretically and observed experimentally; some examples include the thermalization of states far from equilibrium [4, 5, 6, 7, 8, 9, 10] and the description of far-from-equilibrium steady states in terms of state variables that are linked by far-from-equilibrium analogs of equations of state (EoS) [11, 12, 13].
Remarkable examples of steady states that are even locally far from equilibrium are matter-wave turbulent cascades, sustained by energy injection and dissipation that occur on different length scales. Recently, these states have been cleanly realized in a cold-atomic setting [14, 15], and a far-from-equilibrium EoS was measured for the turbulent state [13]. The simplest approach to describe these experiments is the classical-field Gross-Pitaevskii (GP) equation. Although this model has been studied for decades (see e.g. [16, 17]) – and indeed its equilibrium and near-equilibrium properties are well established – it is notable that much of its far-from-equilibrium content remains to be understood  [^1] There also exists a vast body of work on the topic of vortex turbulence in the GP model (see e.g. [19] and references therein).. A successful approach to understanding the GP description in a far-from-equilibrium setting is the perturbative framework of weak wave turbulence (WWT), which predicts the existence of a far-from-equilibrium EoS, the Kolmogorov-Zakharov spectrum [20, 21, 22]. However, this solution does not match the experimentally measured EoS [13], which motivates a non-perturbative study of wave-turbulent cascades in the GP model.
Figure 1: The direct turbulent cascade in the GP model. (a) Cartoon of the simulation geometry and the driving protocol. We use a cylindrical box trapping potential of length $L$ and radius $R$, and the energy is injected into the system by applying a time-periodic potential gradient $V_{\text{drive}}(\bm{r},t)=U_{\textrm{s}}\sin(\omega t)z/L$ (see text for typical parameters). (b) The build-up of the turbulent cascade; the mode occupation number $N_{k}$ is shown for various shaking times $t$. At long times, the system is in a steady state with $N_{k}\propto k^{-\gamma}$ (the dashed line corresponds to $\gamma=3.5$). The inset shows the cascade exponent $\gamma(k)\equiv-\textrm{d}\ln[n(k)]/\textrm{d}k$, calculated from the continuous momentum distribution $n(k)$; the dashed line is $\gamma(k)=3.5$. The red shading indicates the region of momentum space where the weak interaction approximation is not valid (see text). Here, the simulation parameters are $L=50\,\upmu{\rm m}$, $R=15\,\upmu{\rm m}$, $U_{\textrm{s}}=1.0\zeta$, $\omega=2\pi\times 10\,{\rm Hz}$, and $a=100a_{0}$ (corresponding to $\xi=1.2\,\upmu{\rm m}$ and $\tau=10\,$ms). (c) The cascade exponent $\gamma(k)$ for different system parameters. (left) $\gamma(k)$ for different interactions is not universal, and it bends up around $k_{\textrm{D}}/2$ (solid line). (right) The rescaled data for $\gamma$ versus $k\xi$ collapse in the range $22.5\lesssim k\xi\lesssim k_{\textrm{D}}\xi/$ 2, demonstrating that the effective injection scale $k_{0}$ of the isotropic $4$-wave cascade is $\propto k_{\xi}\equiv 1/\xi$. The dashed line shows the theoretical prediction $\gamma(k)-3=1/[3\ln(k/k_{0})]$ with $k_{0}=1.64k_{\xi}$.
In this Letter, we show that the GP model possesses a far-from-equilibrium EoS that lies outside the paradigm of traditional wave-turbulence kinetics [23]. The turbulent state obtained at long times is steady and is characterized by a momentum distribution quantitatively described by a universal prediction from WWT. This steady state possesses the hallmarks of a direct energy cascade, i.e. that the dissipation-scale-independent energy flux is scale invariant, transporting energy from large to small length scales. However, the power-law scaling of the amplitude of the momentum distribution with the energy flux is starkly different from the predictions of wave-kinetics theory.
Our study is based on the universal GP model
$i\hbar\frac{\partial\psi}{\partial t}=\left(-\frac{\hbar^{2}}{2m}\nabla^{2}+g|\psi|^{2}\right)\psi,$ (1)
for the field $\psi(\bm{r},t)$. This equation, also known as the nonlinear Schrödinger equation, is a universal wave equation that describes a variety of systems, such as optical fields in nonlinear Kerr media [24], weakly interacting Bose-Einstein condensates [25], and gravity waves in deep inviscid fluids [26]. We focus on the system of the weakly interacting Bose gas in 3D, where $\psi(\bm{r},t)$ is interpreted as the classical field of the Bose gas, $g=4\pi\hbar^{2}a/m$ is the strength of the interatomic interactions, $m$ is the atomic mass, and $a$ is the s-wave scattering length. The field is normalized to $\int|\psi(\bm{r},t)|^{2}\text{d}^{3}\bm{r}=N$, where $N$ is the (instantaneous) particle number.
Recent experimental studies have shown that this classical field model is relevant to describe a wave-turbulent scale-invariant steady state of a quantum degenerate Bose gas driven at large length scales and dissipated at small ones [14, 15]. To study far-from-equilibrium properties of the GP model Eq. (1), we add trapping, forcing, and dissipation terms in its right-hand side, of the form $V(\bm{r},t)=V_{\text{box}}(\bm{r})+V_{\text{drive}}(\bm{r},t)+V_{\text{diss}}(\bm{r})$ (see [27]). The first contribution, $V_{\text{box}}$, is the confining potential that describes an atomic box trap [28]. For experimental relevance, we use a cylindrical potential (oriented along the $z$ axis) of radius $R$ and length $L$ [see the cartoon of Fig. 1(a)]  [^2] Note that even though the trap shape affects large-scale properties, the far-from-equilibrium state at smaller scales is largely insensitive to the details of the confining potential.. The second term is a forcing at the length scale of the system size, $V_{\text{drive}}(\bm{r},t)=U_{\textrm{s}}\sin(\omega t)z/L$. The natural scale for the drive amplitude $U_{\textrm{s}}$ is set by $gn$, where $n=N/V$ is the average density and $V$ is the trap volume. The third term implements small-length-scale dissipation. We choose $V_{\text{diss}}$ to mimic the dissipation encountered in experiments, i.e. evaporative losses when the atom energy exceeds the trap depth $U_{\textrm{D}}$ [15]. This dissipation term is critical for realizing a steady state in a continuously forced system.
[…]
As shown in Fig. 1(b), the injection of energy results in a cascade front propagating to higher momenta, until hitting the wavenumber $k_{\textrm{D}}$ at time $\approx t_{\textrm{D}}$  [^3] In [27], we derive an analytic expression for $t_{\textrm{D}}$ in terms of the cascade properties.. For $t\gtrsim t_{\textrm{D}}$, the system is in a steady state well described by a power law distribution for the mode occupation number: $N_{k}\propto k^{-\gamma}$ with $\gamma\approx 3.5$. Here, $N_{k}$ is the mode occupation number for the states of momentum $k$, normalized as $\sum_{\bm{k}}N_{k}=N$ (in the language of WWT, $N_{k}$ is the waveaction spectrum); in the continuous limit, it is related to the momentum distribution $n(k)$ as $N_{k}=((2\pi)^{3}/V)n(k)$.
As the momentum $k_{\xi}\equiv 1/\xi$ marks the typical momentum scale associated with interactions, we have that for $k\gg k_{\xi}$, our system is weakly nonlinear, with $4$-wave interactions between (bare) particles; our system can then be described by the theory of WWT [23, 32]. This theory predicts that, in the limit of weak nonlinearities, the momentum distribution of the direct energy cascade has the asymptotic isotropic form $n(k)\propto k^{-3}\ln^{-1/3}(k/k_{0})$ where $k_{0}$ is the energy injection scale [22]. At lower momenta ($k\lesssim k_{\xi}$) the particles interact more strongly, and the $4$-wave WWT description is not expected to work  [^4] In the presence of a condensate, the proper quasiparticles to consider for $k\xi\lesssim 1$ are the Bogoliubov phonons, for which the appropriate WWT description is a wave-kinetic equation with 3-wave interactions. In our simulations, a description in terms of an equilibrium-like condensate is meaningless since for our strong drives, the Bogoliubov approximation is invalid..
Figure 2: Energetics of the direct turbulent cascade. (a) Energy input and dissipation rates. We show $\tau\epsilon/(n\zeta)$ where $\epsilon$ is either the energy injection rate calculated as $\epsilon_{\textrm{in}}\equiv\langle Fvn\rangle$ (symbols) or the particle dissipation rate $\dot{N}$ multiplied by $U_{\textrm{D}}/V$ (solid lines). Both $\epsilon_{\textrm{in}}/n$ and $\dot{N}U_{\textrm{D}}/V$ are constant at long times (see also [27]), but $\epsilon_{\textrm{in}}/n$ is higher by a factor of $\approx 1.3$. The vertical solid lines mark the onset time of dissipation $t_{\textrm{D}}$; for an analytical calculation of $t_{\textrm{D}}$, see [27]. (b) The ratio $\alpha$ of $\epsilon_{\textrm{in}}$ and ${\dot{N}U_{\textrm{D}}/V}$ for different drive strengths $U_{\textrm{s}}/\zeta$ and scattering lengths $a/a_{0}$. Inset: the dissipation spectrum $D_{\varepsilon}$ for $U_{\textrm{s}}=\zeta$ and $a=100a_{0}$. The average dissipation momentum $\langle k_{\textrm{diss}}\rangle\approx 1.15k_{\textrm{D}}$, predicting $\alpha\approx 1.32$ (see text). (c) Energy flux $\Pi_{\varepsilon}$ (top) and particle flux $\Pi_{N}$ (bottom) for different dissipation scales $k_{\textrm{D}}$ (vertical dashed lines). Both fluxes are scale independent. The dotted line ($\propto k^{-2}$) shows that $\Pi_{N}\propto k_{\textrm{D}}^{-2}$, while $\Pi_{\varepsilon}$ is (nearly) independent of $k_{\textrm{D}}$; horizontal dashed lines are $\epsilon_{\textrm{in}}$ (resp. $\tau\dot{N}/N$) in the upper (resp. lower) panel (see text).
However, in any real system, the momentum range over which universal turbulent properties (i.e. injection and dissipation independent, boundary condition independent, etc.) may exist is finite. Thus, the validity of asymptotic results is questionable, and the observations might depend on the separation between the injection and dissipation scales. To test the dependence of the steady-state $n(k)$ on those scales, we calculate the apparent cascade exponent $\gamma(k)\equiv-\textrm{d}\ln[n(k)]/\textrm{d}k$ for different $\xi$ and $k_{\textrm{D}}$. As shown in Fig. 1(c), when $\gamma$ is plotted versus $k$ for various $\xi$ (but same $k_{\textrm{D}}$), it decreases for $k\lesssim k_{\textrm{D}}/2$ and depends on $\xi$; for $k\gtrsim k_{\textrm{D}}/2$, all $\gamma(k)$ increase together. However, when plotted against $k\xi$, the results are universal in the intermediate regime $22.5\lesssim k\xi\lesssim k_{\textrm{D}}\xi/$ 2; this shows that, despite the fact that the energy is physically injected at the scale $k_{\textrm{F}}\approx\pi/L$, the effective injection scale for the isotropic $4$-wave cascade is actually $k_{0}\propto k_{\xi}$ ($\gg k_{\textrm{F}}$ [^5] Interestingly, a recent experiment has observed that the momentum distribution of $2$D shaken Bose gases also becomes isotropic at $k\approx k_{\xi}$ [8].. To test the asymptotic form $n(k)\propto k^{-3}\ln^{-1/3}(k/k_{0})$, we fit $\gamma(k)$ with $3+1/(3\ln^{-1/3}[Ak\xi])$ in the regime where the numerical data overlap. We find that for $A=0.61(1)$, the fit captures the data well. Equivalently, this fit predicts that the injection scale for the isotropic $4$-wave cascade is $k_{0}=1.64(2)k_{\xi}$ (as long as $k_{\textrm{F}}\ll k_{\xi}$  [^6] Experimentally, $k_{\textrm{D}}$ is bounded, so the separation between the injection scale $k_{0}\propto k_{\xi}$ and $k_{\textrm{D}}$ – and hence the region where the $4$-wave WWT theory is applicable – is reduced for stronger interactions.). The convergence towards a robust steady state, i.e. that is independent of $\xi$, $\omega$, $k_{\textrm{D}}$, and $U_{\textrm{s}}$, and that matches a universal expectation that is independent of both the driving protocol and the macroscopic nature of the dissipation, strongly suggests that we have access here to intrinsic properties of the universal GP model Eq. (1).
As was already presumed experimentally [14, 13], we find that a power law $n(k)\propto k^{-\gamma_{0}}$ with an effective exponent $\gamma_{0}=3.5$ accurately captures all the relevant details, since the variation of the logarithmic correction for $\gamma(k)$ is small over the relevant momentum range (see also a discussion in  [27]). From now on, we fix $n(k)$ to this power law and define the amplitude of the power law $n_{0}\equiv N_{k}k^{3}(k\xi)^{\gamma_{0}-3}$.
Figure 3: (a) A universal equation of state of the GP model. Expressing the state variables $\epsilon$ and $n_{0}$ in the instantaneous natural scales of the GP model ($n$, $\xi_{t}$, $\zeta_{t}$, $\tau_{t}$) collapses data at different $a$ onto a universal curve. The dashed line is a power-law fit to the data that gives $n_{0}/n=29(2)(\tau_{t}\epsilon/[n\zeta_{t}])^{0.67(2)}$. (b) The instantaneous far-from-equilibrium state variables for different $t$ follow the EoS, in analogy with an equilibrium-thermodynamics quasi-static process; the data correspond to $a=100a_{0}$, see the green star in panel (a). The points in panel (a) correspond to shaking times $\approx 2t_{\textrm{D}}$ and the green star in (a) is the same point as the blue star in (b). The inset shows the fraction of particles $N/N_{0}$ left in the system.
We next turn to the calculation of the energy-density input and dissipation rates. As energy is injected into the system by the forcing potential $V_{\text{drive}}(\bm{r},t)=U_{\textrm{s}}\sin(\omega t)z/L$, the energy input rate can be calculated as $\epsilon_{\textrm{in}}\equiv\langle Fvn\rangle$ where $F\equiv-\hat{z}\cdot\nabla V_{\text{drive}}$ (where $\hat{z}$ is the unit vector along $z$), $v$ is the center-of-mass velocity of the gas and the averaging is performed over a drive period. The energy is dissipated at high momenta solely by particles with momenta $k>k_{\textrm{D}}$ leaving the trap. Hence, in previous experimental works [15, 13], the energy dissipation rate was determined as $\dot{N}U_{\textrm{D}}$ where $\dot{N}$ is the particle loss rate. However, as shown in Fig. 2(a) for three different $U_{\textrm{s}}$, the energy-density injection rate $\epsilon_{\textrm{in}}$ is consistently higher than $\dot{N}U_{\textrm{D}}/V$. While $\dot{N}U_{\textrm{D}}/V$ is initially zero and increases smoothly to its final value when the system approaches the steady state, this final value is always lower than $\epsilon_{\textrm{in}}$. We investigate this discrepancy systematically by showing the ratio $\alpha=V\epsilon_{\textrm{in}}/(\dot{N}U_{\textrm{D}})$ for different $U_{\textrm{s}}$ and $a$ in Fig. 2(b). The ratio $\alpha\approx 1.3$ is independent of $U_{\textrm{s}}$  [^7] Even though $\alpha\neq 1$ means that the energy-density flux is not $\dot{N}U_{\textrm{D}}/V$, the fact that $\alpha$ has only a very weak dependence on $a$ ($\propto a^{0.08(1)}$) and is independent of $U_{\textrm{s}}$ suggests that the determination of the energy flux via $\dot{N}U_{\textrm{D}}/V$, as done in [15, 13], will only introduce a small, presumably system-specific, systematic error.. The reason for $\alpha\neq 1$ is the shape of the dissipation spectrum $D_{\varepsilon}(k)$ (see [27] for a formal definition): calculating the energy-density dissipation rate as $\dot{N}U_{\textrm{D}}/V$ assumes that the dissipation happens exactly at energy $U_{\textrm{D}}$. As shown in the inset of Fig. 2(b) for a typical $U_{\textrm{s}}=\zeta$ and $a=100a_{0}$, $D_{\varepsilon}(k)$ has a sharp onset at $k_{\textrm{D}}$, but has a significant tail above $k_{\textrm{D}}$. The average dissipation momentum is $\langle k_{\textrm{diss}}\rangle\equiv\int kD_{\varepsilon}(k)\textrm{d}k/\int D_{\varepsilon}(k)\textrm{d}k\approx 1.15k_{\textrm{D}}$, showing that the energy dissipated from the system per unit time is $\approx 1.32\dot{N}U_{\textrm{D}}$.
Having verified that the energy input (at low $k$) and dissipation (at high $k$) rates are equal, we turn to the direct calculation of the scale-resolved energy flux $\Pi_{\varepsilon}(k)$ and particle flux $\Pi_{N}(k)$ (see [27] for the formal definition of the fluxes). As shown in Fig. 2(c), $\Pi_{\varepsilon}$ is momentum independent in steady state and equal to the energy injection rate $\epsilon_{\textrm{in}}$. The energy injected at low $k$ is transported to higher $k$ by the momentum-independent flux $\Pi_{\varepsilon}(k)$ and dissipated at $\gtrsim k_{\textrm{D}}$. As expected for an energy cascade, $\Pi_{\varepsilon}$ is also (nearly) independent of the dissipation scale $k_{\textrm{D}}$  [^8] Note that the slow decrease of the plateau of $\Pi_{\varepsilon}$ is due to a progressive depletion of the condensate, which results in a (slow) decrease of $\epsilon_{\textrm{in}}$ [corresponding horizontal dashed lines in the upper panel of Fig. 2(c)].. On the other hand, $\Pi_{N}$ is also momentum independent in steady state, but its plateau value decreases as $\propto k_{\textrm{D}}^{-2}$, as expected for particles with a quadratic dispersion relation [see dotted line in the lower panel of Fig. 2(c)]  [^9] Note that it is only in the limit $k_{\textrm{D}}\rightarrow\infty$ that one recovers the expected limits $\Pi_{\varepsilon}\rightarrow\epsilon$ with $\Pi_{N}\rightarrow 0$ for the direct energy cascade [20].. We define $\epsilon$, the scale-invariant energy-density flux, as the plateau value of $\Pi_{\varepsilon}$; in the rest of the paper we compute $\epsilon$ as $\langle Fvn\rangle$.
Finally, we investigate the relation between the two far-from-equilibrium state variables of the system, $\epsilon$ and $n_{0}$. Fig. 3(a) shows $n_{0}$ and $\epsilon$ for different $a$; we express both state variables using the instantaneous natural scales of the GP equation $\xi_{t}$, $\tau_{t}$, and $\zeta_{t}$ defined through the instantaneous density $n$  [^10] Note that $\xi_{t}$ is not directly present in the axes of Fig. 3, but is used in the calculation of $n_{0}=N_{k}k^{-3}(k\xi_{t})^{-0.5}$.. Data for different values of $a$ fall on a (nearly) universal curve, demonstrating that the only relevant scales describing the turbulent cascade are the intrinsic scales of the universal GP model Eq. (1) and all the dependence on the system and drive scales ($L$, $R$, $U_{\textrm{s}}$, $\omega$, $k_{\textrm{D}}$) drops out  [^11] The independence of our EoS with $k_{\textrm{D}}$ cannot rigorously hold as for large enough $k_{\textrm{D}}$ (presumably beyond what we explored), WWT results should be recovered. .
We find that the GP model contains a turbulent EoS that is a power law of the form $n_{0}/n=C(\tau_{t}\epsilon/[n\zeta_{t}])^{b}$, with $C=29(2)$ and $b=0.67(2)$, which is not described by known paradigms of turbulence. In particular, it cannot be explained by $4$-wave processes: the relation between $n_{0}$ and $\epsilon$ would have to be independent of the total density $n$ which is not the case for the EoS constructed here  [^12] The factor of $\xi^{0.5}$ on the $y$-axis is a finite-$k_{\textrm{D}}$ effect, but even in its absence our EoS would depend on $n$; within the GP model, only the WWT prediction of a power-law EoS with an exponent of $1/3$ is independent of $n$.. Furthermore, the power-law exponent $0.67(2)$ is inconsistent with any wave-kinetic description, as a kinetic theory with $\ell$-wave interactions predicts an exponent $b=1/(\ell-1)\leq 0.5$. Interestingly, this exponent for the amplitude of mode occupation is numerically the same as the exponent for the amplitude of the energy spectrum for hydrodynamic turbulence $E_{\text{K41}}\propto\epsilon^{2/3}$ [42]. However, this is likely coincidental. First, these quantities are different; one step further: the energy spectrum associated with our $N_{k}$ is $E(k)\propto k^{0.5}$ for $\gamma=3.5$ (assuming a quadratic dispersion relation), while $E_{\text{K41}}\propto k^{-5/3}$. Secondly, the predictions of K41 are independent of interactions, while our universal EoS strongly depends on them (as the rescaling of $\epsilon$ with $\zeta_{t}/\tau_{t}$ corresponds to a rescaling with $a^{2}$).
Finally, we note that the slow decrease of $N$ due to evaporative losses above $k_{\textrm{D}}$ implements a slow thermodynamic-like process. Indeed, as $N$ decreases, the state variable $\epsilon\propto N$ also decreases, resulting in a slow change of the far-from-equilibrium state (see also [27]). In Fig. 3(b), we show that the simultaneous change of the state variables $\epsilon$ and $n_{0}$ follows the EoS: when rescaled to the instantaneous GP scales, the instantaneous $\epsilon$ and $n_{0}$ ‘slide up’ on our universal EoS line over time, see blue to red shades  [^13] As shown in [27], $n_{0}$ and $\epsilon/n$ are roughly constant over time, resulting in the increase of $n_{0}/n$ and $\tau_{t}\epsilon/(n\zeta_{t})$. This is reminiscent of the concept of thermodynamic quasi-static processes, for which infinitesimal changes of external constraints take a system through a dense succession of equilibrium states [44]. Remarkably, we find that this concept generalizes to states that are even locally far from equilibrium.
In conclusion, we numerically investigated the properties of a turbulent cascade arising in the GP model when a box trapped gas is periodically driven and showed that it can be described by a universal EoS relating the turbulent state variables, the energy flux and the cascade amplitude. The form of our EoS is inconsistent with any kinetic theory, posing a new theoretical challenge. The comparision of our mean-field simulations to experimental measurements provides valuable benchmarks for testing the validity of mean-field theories in far-from-equilibrium scenarios. Note that even though our work provides some elements to understand the experimental results of [13], the experimental EoS cannot be fully understood solely within the framework of the GP model as the data features a scaling with $na^{3}$ that is incompatible with the universal GP model (see also [27]). In the future, it would be interesting to study on its own the incompressible parts of this turbulent scenario [19]. One should also investigate the fluctuations of $N_{k}$; comparing these fluctuations between experiments and classical field simulations could shed light on the role of quantum fluctuations in turbulent quantum fluids.
We thank M. Gazo, J. Etrych, C. Eigen, G. Krstulovic, M. Tsubota, G. Falkovich, and V. L’vov for discussions. This work was supported by the NSF (Grant Nos. PHY-1945324 and PHY-2110303), DARPA (Grant No. HR00112320038), the David and Lucile Packard Foundation, the Alfred P. Sloan Foundation, and JSPS KAKENHI (Grant No. JP23K13029). G.M. acknowledges support from Trinity College (Cambridge) and Yale University for its hospitality.