### v2

# arXiv:2407.08738v2[cond-mat.quant-gas] 28 Feb 2026

Abstract
We report the numerical observation of a far-from-equilibrium equation of state (EOS) in the Gross-Pitaevskii model. We first show that the momentum distribution of the turbulent cascade is well described by wave-turbulent kinetic theory in the appropriate limits. Calculating the energy and particle fluxes $\Pi_{\varepsilon}(k)$ and $\Pi_{N}(k)$, we show that the turbulent state possesses the hallmarks of a direct energy cascade. Building on this, we show that the GP model encodes a universal EOS in the form of a relationship between the turbulent cascade’s momentum distribution amplitude $n_{0}$ and the energy flux $\epsilon$ in the steady state. We find that in our regime of ‘mixed’ turbulence - where both vortices and waves play a significant role - $n_{0}\propto\epsilon^{0.67(2)}$, a result that is not captured by any existing theory of turbulence but that agrees with a recent experimental measurement for large energy fluxes. Finally, we find that the concept of quasi-static thermodynamic processes between equilibrium states extends to far-from-equilibrium steady states.
Introduction. Equilibrium and near-equilibrium thermodynamics form conceptual cornerstones of physics, reducing the behavior of complex many-body systems to relations among a few macroscopic observables. Far from equilibrium, such a unifying framework is lacking, and unveiling universal features has become a major goal of modern physics. When local equilibrium holds, local thermodynamics and hydrodynamics provide a bridge to universal descriptions [1, 2, 3]. Even without microscopic equilibrium, universal phenomena have been predicted and observed—ranging from thermalization of far-from-equilibrium states [4, 5, 6, 7, 8, 9, 10, 11] to steady states characterized by state variables linked through far-from-equilibrium analogs of equations of state (EOS) [12, 13, 14].
A remarkable example of steady states that remain locally far from equilibrium is matter-wave turbulence, sustained by energy injection and dissipation at distinct length scales. Such cascades have recently been realized in ultracold-atom systems [15, 16, 17], where a far-from-equilibrium EOS was measured [14]. All tractable theoretical descriptions of these experiments rely on the classical-field Gross–Pitaevskii (GP) equation [18, 19, 20, 21] [^1] This model is not limited to weakly interacting BECs; it also applies in the classical-field limit of highly occupied modes [54, 55]..
Although the GP model has been studied for decades [23, 24], and its equilibrium and near-equilibrium behavior are well established, much of its far-from-equilibrium physics remains to be understood. It supports regimes such as Kolmogorov vortex turbulence  [25, 26, 27, 28, 29, 17] and weak-wave turbulence (WWT) [30, 31], which feature a far-from-equilibrium EOS relating the amplitude of a steady-state spectrum to an energy flux. However, these solutions do not describe the recently measured experimental EOS [14], motivating a calculation of the EOS within the GP model in the same setting as the experiments.
In this Letter, we perform simulations of the GP model in the setting of the experiments [14] and show that the GP model possess a far-from-equilibrium EOS that lies outside any known paradigm of turbulence. The turbulent state obtained at long times is steady and we find that it is characterized by a momentum distribution quantitatively described by a universal prediction from WWT theory. This steady state possesses the hallmarks of a direct energy cascade, i.e. that the dissipation-scale-independent energy flux is scale invariant, transporting energy from large to small length scales. However, the power-law scaling of the amplitude of the momentum distribution with the energy flux is starkly different from predictions of any theory of turbulence, but agrees well with the experimental EOS for large values of the energy flux.
Figure 1: The direct turbulent cascade in the GP model. (a) Cartoon of the simulation geometry and the driving protocol. We use a cylindrical box trapping potential of length $L$ and radius $R$, and the energy is injected into the system by applying a time-periodic potential gradient $V_{\text{drive}}(\bm{r},t)=U_{\textrm{s}}\sin(\omega t)z/L$ (see text for typical parameters). (b) The build-up of the turbulent cascade; the mode occupation number $N_{k}$ is shown for various shaking times $t$. At long times, the system is in a steady state with $N_{k}\propto k^{-\gamma}$ (the dashed line corresponds to $\gamma=3.5$). The inset shows the cascade exponent $\gamma(k)\equiv-\textrm{d}\ln[n(k)]/\textrm{d}\ln{[k]}$, calculated from the continuous momentum distribution $n(k)$; the dashed line is $\gamma(k)=3.5$. Here, the simulation parameters are $L=50\,\upmu{\rm m}$, $R=15\,\upmu{\rm m}$, $U_{\textrm{s}}=1.0\zeta$, $\omega=2\pi\times 10\,{\rm Hz}$, and $a=100a_{0}$ (corresponding to $\xi=1.2\,\upmu{\rm m}$ and $\tau=10\,$ms).
The model. Our study is based on the universal GP model
$i\hbar\frac{\partial\psi}{\partial t}=\left(-\frac{\hbar^{2}}{2m}\nabla^{2}+g|\psi|^{2}\right)\psi,$ (1)
for the field $\psi(\bm{r},t)$. This equation, also known as the nonlinear Schrödinger equation, is a universal wave equation that describes a variety of systems, such as optical fields in nonlinear Kerr media [32], weakly interacting Bose-Einstein condensates [33], and gravity waves in deep inviscid fluids [34]. We focus on the system of the weakly interacting Bose gas in 3D, where $\psi(\bm{r},t)$ is interpreted as the classical field of the Bose gas, $g=4\pi\hbar^{2}a/m$ is the strength of the interatomic interactions, $m$ is the atomic mass, and $a$ is the s-wave scattering length  [^2] It is worthwhile noting that, as in the recent experiments [15, 14], our study remains in a regime $ka\ll 1$, where $k$ is the highest momentum considered. The scattering cross-section is thus momentum independent and $g=4\pi\hbar^{2}a/m$.. The field is normalized to $\int|\psi(\bm{r},t)|^{2}\text{d}^{3}\bm{r}=N$, where $N$ is the (instantaneous) particle number.
To study universal steady-state properties of the GP model far from equilibrium, one has to supplement forcing and dissipation mechanisms into Eq. (1); here we do that by adding a potential term of the form $V(\bm{r},t)=V_{\text{drive}}(\bm{r},t)+V_{\text{diss}}(\bm{r})+V_{\text{box}}(\bm{r})$ (see [36]). The first term is a forcing at the length scale of the system size $L$, $V_{\text{drive}}(\bm{r},t)=U_{\textrm{s}}\sin(\omega t)z/L$. The second term implements small-length-scale dissipation. This dissipation term is critical for realizing a steady state in a continuously forced system; we choose $V_{\text{diss}}$ to mimic the dissipation encountered in experiments, i.e. evaporative losses when the atom energy exceeds an energy $U_{\textrm{D}}$ [16]. The last term, $V_{\text{box}}$, is a confining potential; for experimental relevance we pick a cylindrical box potential [37, 14] whose axis is oriented along the shaking direction $z$ [see the cartoon of Fig. 1(a)]  [^3] Note that even though the trap shape affects large-scale properties, the far-from-equilibrium state at smaller scales is largely insensitive to the details of the confining potential.. Recent experimental studies have confirmed that this classical field approach is useful for describing a wave-turbulent scale-invariant steady state of a quantum degenerate Bose gas [15, 16].
The system is initialized in the ground state of Eq. (1) including $V_{\text{box}}(\bm{r})$, which corresponds to a Bose-Einstein condensate with a nearly uniform density, except near the boundary of the box. For relevance, we use numbers typical of recent experiments [14]: a box radius $R=15\,\upmu{\rm m}$ and length $L=50\,\upmu{\rm m}$, an initial atom number $N(t=0)=2\times 10^{5}$, and $a/a_{0}$ between $25$ and $400$ (where $a_{0}$ is the Bohr radius). The natural energy scale of the system $\zeta\equiv gn(t=0)$ therefore varies between $k_{\textrm{B}}\times 1.2\,{\rm nK}$ and $k_{\textrm{B}}\times 19\,{\rm nK}$, the corresponding timescale $\tau\equiv h/\zeta$, between $2.6\,\textrm{ms}$ and $42\,\textrm{ms}$, and the natural lengthscale $\xi\equiv\sqrt{2m\zeta}/\hbar$, between $0.6\,\upmu{\rm m}$ and $2.3\,\upmu{\rm m}$ (using the mass $m$ of 39K); $n=N/V$ is the average density, where $V=\pi R^{2}L$ is the box volume. The gas is driven at a frequency $\omega/(2\pi)$ that matches the resonance of the lowest-lying Bogoliubov excitation [39]; for our set of parameters, $\omega/(2\pi)$ ranges from $5$ to $20\,\textrm{Hz}$.
Momentum distribution. As shown in Fig. 1(b), the injection of energy results in a cascade front propagating to higher momenta, until hitting the momentum $k_{\textrm{D}}$  [^4] For convenience, we refer to the wavenumber k as momentum, even though formally the momentum is $\textbf{p}=\hbar\textbf{k}$. at a time $t\approx t_{\textrm{D}}$  [^5] In End Matter [36], we derive an analytic expression for $t_{\textrm{D}}$ in terms of the cascade properties.. For $t\gtrsim t_{\textrm{D}}$, the system is in a steady state well described by a power law distribution for the mode occupation number: $N_{k}\propto k^{-\gamma}$ with $\gamma\approx 3.5$. Here, $N_{k}$ is the mode occupation number for the states of momentum $k$, normalized as $\sum_{\bm{k}}N_{k}=N$; in the continuous limit, it is related to the momentum distribution $n(k)$ as $N_{k}=((2\pi)^{3}/V)n(k)$.
We systematically studied the momentum-resolved cascade exponent $\gamma(k)\equiv-\textrm{d}\ln[n(k)]/\textrm{d}\ln{[k]}$ across interaction strengths (different $\xi$) and find that $\gamma(k\xi)$ is in excellent agreement with the $4$-wave WWT prediction $3+1/(3\ln^{1/3}[k/k_{\textrm{F}}])$, provided that $k_{\textrm{F}}$ is replaced by $k_{0}=1.64(2)k_{\xi}$ (see End Matter and Fig. 5). In other words, the isotropic $4$-wave cascade’s effective injection scale is set by the GPE scale $1/\xi$ instead of the physical injeciton scale $k_{\textrm{F}}$. However, as was already presumed experimentally [15, 14], we find that a power law $n(k)\propto k^{-\gamma_{0}}$ with an effective exponent $\gamma_{0}=3.5$ accurately captures all the relevant details, since the variation of the logarithmic correction for $\gamma(k)$ is small over the relevant momentum range  [^6] Note that [20] speculated that the steeper effective exponent could be due to the logarithmic correction. However, this work assumes that there is no BEC, but the presence of the BEC is crucial for understanding the effective injection scale $k_{0}\propto k_{\xi}$, hence [20] cannot make quantitative comparisons to the experiment [15]. (see also a discussion in  [36]). From now on, we fix $n(k)$ to this power law and define the amplitude of the power law $n_{0}\equiv N_{k}k^{3}(k\xi)^{\gamma_{0}-3}$.
Figure 2: Energetics of the direct turbulent cascade. (a) Energy input and dissipation rates. We show $\tau\epsilon/(n\zeta)$ where $\epsilon$ is either the energy injection rate calculated as $\epsilon_{\textrm{in}}\equiv\langle Fvn\rangle$ (symbols) or the particle dissipation rate $\dot{N}$ multiplied by $U_{\textrm{D}}/V$ (solid lines). Both $\epsilon_{\textrm{in}}/n$ and $\dot{N}U_{\textrm{D}}/N$ are constant at long times (dashed lines; see also [36]), but $\epsilon_{\textrm{in}}/n$ is higher by a factor of $\approx 1.3$. The vertical solid lines mark the onset time of dissipation $t_{\textrm{D}}$; for an analytical calculation of $t_{\textrm{D}}$, see the End Matter. Inset: the dissipation spectrum $D_{\varepsilon}$ for $U_{\textrm{s}}=\zeta$ and $a=100a_{0}$. The average dissipation momentum $\langle k_{\textrm{diss}}\rangle\approx 1.15k_{\textrm{D}}$, predicting $\epsilon V/(\dot{N}U_{\textrm{D}})\approx 1.32$ (see text). (b) Energy flux $\Pi_{\varepsilon}$ (left) and particle flux $\Pi_{N}$ (right) for different dissipation scales $k_{\textrm{D}}$ (vertical dashed lines). Both fluxes are scale independent. The dotted line ($\propto k^{-2}$) shows that $\Pi_{N}\propto k_{\textrm{D}}^{-2}$, while $\Pi_{\varepsilon}$ is (nearly) independent of $k_{\textrm{D}}$; horizontal dashed lines are $\epsilon_{\textrm{in}}$ (resp. $\tau\dot{N}/N$) in the left (resp. right) panel (see text).
Energetics of the turbulent cascade. We next turn to the calculation of the energy-density input and dissipation rates. As energy is injected into the system only by the forcing $V_{\text{drive}}$, the energy input rate can be calculated as $\epsilon_{\textrm{in}}\equiv\langle Fvn\rangle$ where $F\equiv-\hat{z}\cdot\nabla V_{\text{drive}}$, $\hat{z}$ is the unit vector along $z$, $v$ is the center-of-mass velocity of the gas and the averaging $\langle\cdot\rangle$ is performed over a drive period. The energy is dissipated at high momenta solely by particles with momenta $k>k_{\textrm{D}}$ leaving the trap. Hence, in previous experimental works [16, 14], the energy dissipation rate was reasonably assumed to be $\dot{N}U_{\textrm{D}}$ where $\dot{N}$ is the particle loss rate. However, as shown in Fig. 2(a) for three different $U_{\textrm{s}}$, the energy-density injection rate $\epsilon_{\textrm{in}}$ is consistently higher than $\dot{N}U_{\textrm{D}}/V$. The reason for $\epsilon_{\textrm{in}}>\dot{N}U_{\textrm{D}}/V$ stems from the shape of the dissipation spectrum $D_{\varepsilon}(k)$ (see [36] for a formal definition): calculating the energy-density dissipation rate as $\dot{N}U_{\textrm{D}}/V$ assumes that the dissipation happens exactly at energy $U_{\textrm{D}}$. As shown in the inset of Fig. 2(a) for $U_{\textrm{s}}=\zeta$, $D_{\varepsilon}(k)$ has a sharp onset at $k_{\textrm{D}}$, but has a significant tail above $k_{\textrm{D}}$. The average dissipation momentum is $\langle k_{\textrm{diss}}\rangle\equiv\int kD_{\varepsilon}(k)\textrm{d}k/\int D_{\varepsilon}(k)\textrm{d}k\approx 1.15k_{\textrm{D}}$  [^7] A simple geometric picture is that particles hitting the box at non-normal incidence can remain trapped for some time even if their kinetic energy $\hbar^{2}k^{2}/(2m)>U_{\textrm{D}}$., showing that the energy dissipated from the system per unit time is $\approx 1.32\dot{N}U_{\textrm{D}}$. This is in excellent agreement with $\alpha\equiv\epsilon_{\textrm{in}}/(\dot{N}U_{\textrm{D}})$  [^8] Even though $\alpha\neq 1$ means that the energy-density flux is not $\dot{N}U_{\textrm{D}}/V$, the fact that $\alpha$ has only a very weak dependence on $a$ ($\propto a^{0.08(1)}$) and is independent of $U_{\textrm{s}}$ suggests that the determination of the energy flux via $\dot{N}U_{\textrm{D}}/V$, as done in [16, 14], will only introduce a small, presumably system-specific, systematic error. [see the End Matter for a systematic study of $\alpha$ versus interaction strength and $U_{\textrm{s}}$, showing that $\alpha\approx 1.3$].
Having verified that the energy input rate (at low $k$) and dissipation rate (at high $k$) are equal, we turn to the direct calculation of the scale-resolved energy flux $\Pi_{\varepsilon}(k)$ and particle flux $\Pi_{N}(k)$ (see [36] for the formal definition of the fluxes). As shown in Fig. 2(c) in the steady state, $\Pi_{\varepsilon}$ is momentum independent in the inertial range, a certain range of momenta where neither forcing nor dissipation takes place; this flux transports the energy injected at low $k$ to higher $k$, up to $k_{\textrm{D}}$. Importantly, $\Pi_{\varepsilon}$ in the inertial range is equal to $\epsilon_{\textrm{in}}$. As expected for an energy cascade, $\Pi_{\varepsilon}$ is also (nearly) independent of the dissipation scale $k_{\textrm{D}}$  [^9] Note that the slow decrease of the plateau of $\Pi_{\varepsilon}$ is due to a progressive depletion of the condensate, which results in a (slow) decrease of $\epsilon_{\textrm{in}}$ [corresponding horizontal dashed lines in the upper panel of Fig. 2(c)].. On the other hand, $\Pi_{N}$ is also momentum independent in steady state, but its plateau value decreases as $\propto k_{\textrm{D}}^{-2}$, as expected for particles with a quadratic dispersion relation [see dotted line in the lower panel of Fig. 2(c)]  [^10] Note that it is only in the limit $k_{\textrm{D}}\rightarrow\infty$ that one recovers the expected limits $\Pi_{\varepsilon}\rightarrow\epsilon$ with $\Pi_{N}\rightarrow 0$ for the direct energy cascade [56].. We define $\epsilon$, the scale-invariant energy-density flux, as the plateau value of $\Pi_{\varepsilon}$; in the rest of the paper we compute $\epsilon$ as $\langle Fvn\rangle$.
Figure 3: A universal equation of state of the GP model. (a) The numerical calculations for different $a$ collapse onto a universal curve when the state variables $\epsilon$ and $n_{0}$ are expressed in the instantaneous natural scales of the GP model ($n$, $\xi_{t}$, $\zeta_{t}$, $\tau_{t}$). The dashed line is a power-law fit to the data that gives $n_{0}/n=29(2)(\tau_{t}\epsilon/[n\zeta_{t}])^{0.67(2)}$, and the pink band shows the fit uncertainty. (b) The comparison of our numerical EOS (gray symbols) with the experimental data from [14] (colored symbols). The color coding of the experimental data is based on the gas parameter $na^{3}$; note that the experimental fluxes are multiplied by $1.3$ compared to the data of [14] to account for $\alpha\neq 1$. The vertical error bars of the experimental data represent the uncertainties due to different $\gamma_{0}$ in simulations and experiments (see [36]).
EOS for turbulence. Finally, we investigate the relation between the two far-from-equilibrium state variables of the system, $\epsilon$ and $n_{0}$. Fig. 3(a) shows $n_{0}$ and $\epsilon$ for different $a$; we express both state variables using the instantaneous natural scales of the GP equation $\xi_{t}$, $\tau_{t}$, and $\zeta_{t}$ defined through the instantaneous density $n$  [^11] Note that $\xi_{t}$ is not directly present in the axes of Fig. 3, but is used in the calculation of $n_{0}=N_{k}k^{-3}(k\xi_{t})^{-0.5}$.. Data for different values of $a$ fall on a (nearly) universal curve, demonstrating that the only relevant scales describing the turbulent cascade are the intrinsic scales of the universal GP model Eq. (1) and all the dependence on the system and drive scales ($L$, $R$, $U_{\textrm{s}}$, $\omega$, $k_{\textrm{D}}$) drops out  [^12] The independence of our EOS with $k_{\textrm{D}}$ presumably will not hold for large enough $k_{\textrm{D}}$ (beyond what we studied), as it is reasonable to expect that as $k\rightarrow\infty$ (with $k<k_{\textrm{D}}$), WWT results should be recovered..
The data of Fig. 3(a) show that the GP model contains a turbulent EOS that is a power law of the form $n_{0}/n=C(\tau_{t}\epsilon/[n\zeta_{t}])^{b}$, with $C=29(2)$ and $b=0.67(2)$, which is not described by known paradigms of turbulence. The two broad paradigms are wave (compressible-energy dominated) and vortex (incompressible-energy dominated) turbulence. Our $b$ is inconsistent with any wave-kinetic description, as a kinetic theory with $\ell$-wave interactions predicts an exponent $b=1/(\ell-1)\leq 0.5$. Furthermore, the most plausible $4$-wave theory predicts that the relation between $n_{0}$ and $\epsilon$ would have to be independent of the total density $n$, which is not the case for the EOS constructed here  [^13] For an extensive comparison of our EOS to the prediction of $4$-wave theory (as well as how the respective $N_{k}$ can be deduced), see Section VII and Fig. S3141414The factor of $\xi^{0.5}$ on the $y$-axis is a finite-$k_{\textrm{D}}$ effect, but even in its absence our EOS would depend on $n$; within the GP model, only the WWT prediction of a power-law EOS with an exponent of $1/3$ is independent of $n$..
To go further, we decompose in [36] the energy spectrum into compressible and incompressible parts [25] and show that while the compressible part is typically larger than the incompressible one in the $k$ range used to extract $n_{0}$, the two are of the same magnitude; moreover, the incompressible part increases more rapidly with $\epsilon$ (similar observations were also recently reported in [21]). The observed EOS is therefore likely a result of an interplay between compressible (wave) and incompressible (vortex) excitations. As far as we are aware, there is no existing theory that describes this regime of ‘mixed’ quantum turbulence and the shape of the EOS remains to be understood. We further note that the value of $b$ being the same as the exponent of the scaling of the energy spectrum for vortex (i.e. hydrodynamic) turbulence $E_{\text{K41}}\propto\epsilon^{2/3}$ [51] is likely coincidental, as this prediction is for the incompressible energy spectrum while our EOS is for the momentum distribution; besides, our computed incompressible energy spectra are also distinct from the K4141 prediction [36]. We note, however, that Kolmogorov turbulence can also be realized and measured in ultracold-atom experiments [17] (for $k\xi\ll 1$) but in that case the K4141 spectrum originates from (coarse-grained) velocity field rather than the (atomic) momentum distribution.
Figure 4: Quasi-static process far from equilibrium. The instantaneous far-from-equilibrium state variables for different $t$ follow the EOS, in analogy with a quasi-static process in equilibrium thermodynamics. The data correspond to $a=100a_{0}$ and the blue star here corresponds to the green star in Fig. 3(a). The inset shows the fraction of particles $N/N_{0}$ left in the system.
In Fig. 3(b) we compare our calculations (grey symbols) with the experimental measurements [14] (colored symbols); see also [36] for details. The experimental data are above the numerical results. Note that the experimental data do not represent a universal EOS on this plot: they collapse onto a single curve when both axes are rescaled with empirically determined powers of the gas parameter $na^{3}$ [14], a scaling that is inherently beyond the GP model. However, we note that in the limit $n\rightarrow\infty$ and $na^{3}\rightarrow 0$ the GP model is expected to be a good description of the experimental setting and Fig. 3(b) shows that the experimental data appear to approach our numerical results for lower $na^{3}$.
Finally, we note that the slow decrease of $N$ due to evaporative losses above $k_{\textrm{D}}$ implements a slow thermodynamic-like process. Indeed, as $N$ decreases, the state variable $\epsilon\propto N$ also decreases, resulting in a slow change of the far-from-equilibrium state (see also [36]). In Fig. 4, we show that the simultaneous change of the state variables $\epsilon$ and $n_{0}$ follows the EOS: when rescaled to the instantaneous GP scales, the instantaneous $\epsilon$ and $n_{0}$ ‘slide up’ on our universal EOS line over time, see blue to red shades  [^15] As shown in [36], $n_{0}$ and $\epsilon/n$ are roughly constant over time, resulting in the increase of $n_{0}/n$ and $\tau_{t}\epsilon/(n\zeta_{t})$. The points in Fig. 3(a) correspond to shaking times $\approx 2t_{\textrm{D}}$.. This is reminiscent of the concept of thermodynamic quasi-static processes, for which infinitesimal changes of external constraints take a system through a dense succession of equilibrium states [53]. Remarkably, we find that this concept generalizes to states that are even locally far from equilibrium.
Conclusions. We numerically investigated the properties of a turbulent cascade arising in the GP model when a box trapped gas is periodically driven, and showed that it can be described by a universal EOS relating the turbulent state variables, the energy flux and the cascade amplitude. The form of our EOS is inconsistent with any existing theory of turbulence, posing a new theoretical challenge. Furthermore, the comparison of our classical-field simulations to experimental measurements provides valuable benchmarks for testing the validity of classical-field theories in far-from-equilibrium scenarios. Namely, we find that the GP model provides valuable insight into the experimental results of [14], even if it fails to fully capture the equation of state – indeed, the observed $na^{3}$ scaling is incompatible with the universal GP framework (see also [36]). This partial success is nevertheless striking given that there is no firm theoretical justification for its applicability in this turbulent regime. Understanding the foundations of this success, and developing refined approaches that incorporate quantum effects, remain important challenges for the field. One should also investigate the fluctuations of $N_{k}$; comparing these fluctuations between experiments and classical field simulations could shed light on the role of quantum fluctuations in turbulent quantum fluids.
We thank M. Gazo, J. Etrych, C. Eigen, G. Krstulovic, M. Tsubota, G. Falkovich, and V. L’vov for discussions. This work was supported by the NSF (Grant Nos. PHY-1945324 and PHY-2110303), DARPA (Grant No. HR00112320038), the David and Lucile Packard Foundation, the Alfred P. Sloan Foundation, and JSPS KAKENHI (Grant No. JP23K13029). G.M. acknowledges support from Trinity College (Cambridge) and Yale University for its hospitality.